Determine the roots of the equation x^2+4x=45 can you help me in the first steps
Answer by Cromlix(4381)

x^2 + 4x = 45
Change the side of the 45
changing its sign
x^2 + 4x - 45= 0
Now you need to find two numbers
that multiply together to give you -45
and add together to give you +4
5 x 9 = 45
Keep the larger number positive
and make the smaller on negative
-5 x +9 = -45
-5 + 9 = 4
(x - 5)(x + 9) = 0
Hope this helps.
:-)